Complete the following chart Patient Cockcroft –Gault CKD-EPI eGFR BSA Un-normalized CKD-EPI eGFR Patient A 45...

  1. Complete the following chart

Patient

Cockcroft –Gault

CKD-EPI eGFR

BSA

Un-normalized CKD-EPI eGFR

Patient A

45 y.o. white male

SCr 2 mg/dL

5’8” 50 kg

Patient B

45 y.o. white male

SCr 2 mg/dL

5’8” 75 kg

Patient C

45 y.o. white male

SCr 2 mg/dL

5’8” 100 kg
Answer #1

1. Patient A, 45 years old , white male , SCr-2 mg/dl , height - 172.2cm , weight -50kg

Cockcroft gault formula - (140-age ) × body weight ÷ serum creatinine clearance × 72

= (140-45) × 50 / 2 × 72

= 95×50/ 2×72

= 32.8

BSA = √ weight in kg × height in cm /3600

= √ 50× 172.2 /3600

= 1.5m square

2. Patient B ,45years , white male , serum creatinine 2 mg/dl , weight -75kg ,height - 172.2cm

Cockcroft gault formula - (140-age ) × body weight ÷ serum creatinine clearance × 72

= (140-45) × 75/ 2×72

= 49.4

BSA - √ weight in kg × height in cm /3600

= √ 75× 172.2/3600

= 1.8m square

3. atient A, 45 years old , white male , SCr-2 mg/dl , height - 172.2cm , weight -100kg

Cockcroft gault formula - (140-age ) × body weight ÷ serum creatinine clearance × 72

= (140-45) × 100/ 2 × 72

= 65.9

BSA - √ weight in kg × height in cm /3600

= √ 100× 172.2/3600

= 2.1m square